OpenStudy (anonymous):

9(x - 6) = 8(x + 6)?

OpenStudy (anonymous):

9x-54=8x+48

OpenStudy (anonymous):

i am distributing but I forget the next step

OpenStudy (anonymous):

what is 9x-8x

OpenStudy (anonymous):

it is just x

OpenStudy (anonymous):

okay so then you have x-54=48 how do you get that -54 over to the 48

OpenStudy (anonymous):

so x=-10

OpenStudy (anonymous):

no add 54 to 48

OpenStudy (anonymous):

to get x by itself

OpenStudy (anonymous):

so x=102?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now check your work

OpenStudy (anonymous):

9(x - 6) = 8(x + 6) 9x-54=8x+48 9x-4x=x=54+48=102